How do you substitute .05m+20+20=.10(m-100)+15+20

Answer by stanbon(75887)

0.05m+20+20 = 0.10(m-100)+15+20
--------
0.05m + 40 = 0.10m-10 + 35
-----
0.05m + 40 = 0.10m+25
------
0.05m = 15
----
m = 1500/5
----
m = 300
==============
Cheers,
Stan H.